# Spins of Black Holes in X-ray Binaries and the Tension with the Gravitational Wave Measurements

### v2

@@ 3.3 Jet power and spins
Transient ejecta achieve the highest power among jets launched by BH XRBs during hard-to-soft state transitions, with a prominent example of MAXI J1348–630 (Carotenuto et al., 2021). The jet power has been estimated based on the kinetic energy obtained from modeling the propagation of those jets (observed in radio) up to distances up to $\sim$1 pc through the surrounding media together with the jet launching times assumed to equal the durations of major radio flares accompanying the ejecta. It is comparable to the maximum power of Equation (8) at $a_{*}\sim 1$ in several cases (Carotenuto et al., 2024; Cooper et al., 2025). The vast propagation distances require the presence of surrounding cavities with the densities $\ll 1$ cm-3 (Heinz, 2002; Sikora and Zdziarski, 2023). Notably, such powerful ejections have been observed only among LMXBs. Thus, some of them may be required to have high spins. On the other hand, Fender and Motta (2025) have compared the measured ejecta velocities with the published values of the spin obtained using the reflection, continuum, and QPO methods, and found no correlations. This shows that either the ejecta power is at most weakly dependent on the spin, or the published spin values are unreliable (which we advocate here), or both.
@@ 3.4 X-ray polarization and spins
An independent method for measuring the BH spin is through X-ray polarization. The image of the innermost regions of the accretion disk is distorted by general and special relativity effects, causing different parts of the disk – whose emission dominates at different energies – to have varying orientations. This effect is expected to produce a gradual change of PA as a function of energy and progressive decrease of PD (Dovčiak et al., 2004a; Loktev et al., 2022, 2024). The rate of PA rotation and the suppression of PD are sensitive to BH spin, making them valuable diagnostic tools. However, X-ray polarization measurements of soft-state sources show no statistically significant evidence of PA dependence on energy, nor do they indicate a decrease of PD with increasing energy (Svoboda et al., 2024; Marra et al., 2024; Ratheesh et al., 2024; Steiner et al., 2024).
The importance of returning radiation was put forward to explain the observed increase of PD with energy and the nearly constant PA in two sources, 4U 1915+1054 and Cyg X-1 (Marra et al., 2024; Steiner et al., 2024). However, satisfying both criteria required an assumption of an extreme albedo of unity. Moreover, no simultaneous fits of both the spectral and polarimetric signatures were done for the latter: either spectral modelling was performed separately from polarimetric modeling, or the parameters that provided a good fit for the polarization data failed to align with the spectral constraints. For the former, the albedo was found to be not constrained.
@@ 3.5 Published values of the BH spins of XRBs
We know three BH HMXBs with measured spins: Cyg X-1, LMC X-1, and M33 X-7. Their spin values published before 2024 were all high: >>0.9985 (Zhao et al., 2021b; Miller-Jones et al., 2021), $0.92^{+0.05}_{-0.07}$ (Gou et al., 2009), and $0.84\pm 0.05$ (Liu et al., 2008), respectively. Those results were obtained using the disk continuum method in the soft state, specifically with kerrbb2. A new value for Cyg X-1 was found in the spectral fit with kerrbb by Steiner et al. (2024), $a_{*}=0.99964^{+0.00003}_{-0.00007}$. Note that both of the values for Cyg X-1 are above the limit of Thorne (1974), and even more that of Mummery (2025) ($a_{*{\rm max}}\lesssim 0.99$). Although those errors do not include systematic errors and uncertainties in the mass, distance, and inclination, very high values of the spin of Cyg X-1 were also reported using the previous determination of the mass and distance. In particular, Gou et al. (2014) found $a_{*}>0.983$ using the values of $M=14.8{M}_{\odot}$ and $D=1.86$ kpc (Orosz et al., 2011).
Similarly, the published spin values for BH LMXBs are high on average, with $a_{*}\gtrsim 0.9$ in many cases and the mean value of $\langle a_{*}\rangle>0.7$; see the compilations in Reynolds (2021) and Fishbach and Kalogera (2022). A recent paper by Draghis et al. (2024) used the reflection method for a large sample of BH XRBs and claimed that the spins obtained in their work are even higher for most sources than those found previously. Namely, about 86% of their sample had spins consistent with $a_{*}\geq 0.95$, 94% were consistent with $a_{*}\geq 0.9$, and 100% of their spin values were consistent with $a_{*}>0.7$. However, they assumed that the disk extends down to the ISCO in all cases, in both the hard and soft states. In the former, this assumption is highly uncertain (or even unlikely); see Section 3.1.1. Taking their results at face value, this would imply even larger spins, with the obtained values being the lower limits. On the other hand, Draghis et al. (2024) fitted models with the incident spectrum being either an e-folded power law or a single thermal Comptonization spectrum. As we argued in Section 3.1.1, the actual incident spectra are significantly more complex.
@@
Then, the spins have been measured to be low in some binaries while they showed powerful jets, apparently requiring high spins to power them (Equation 8). A prominent case is that of MAXI J1820+070. Using the continuum method, Guan et al. (2021) and Zhao et al. (2021a) obtained $a_{*}=0.2^{+0.2}_{-0.3}$ and $a_{*}=0.14\pm 0.09$, respectively, while Fabian et al. (2020) argued for $a_{*}<0$ based on the possible presence of emission from the plunging region below the ISCO (see Section 3.1.2). On the other hand, this system showed a powerful transient jet (Bright et al., 2020; Carotenuto et al., 2024), which requires $a_{*}\sim 1$ (Zdziarski and Heinz, 2024).
@@ 4.1 Revised spin measurements and improved disk models
Figure 6: A soft-state state of Cyg X-1 observed by NICER, NuSTAR and INTEGRAL (based on Zdziarski et al. 2024b). We show the unfolded spectra and data-to-model ratios in the top and bottom panels. In (a), the spectrum is fitted by the disk model kerrbb and Comptonization yielding $a_{*}=0.87^{+0.04}_{-0.03}$. In (b), the spectrum is fitted with a similar model but including a warm corona, and $a_{*}=0.00^{+0.07}$. The total model spectra and the unabsorbed ones are shown by the solid blue and blue curves, respectively. The unabsorbed disk, scattered, and reflected components are shown by the magenta, green, and red curves, respectively. In (b), the brown curve represents the underlying disk spectrum before it undergoes Comptonization in the top layer. See Zdziarski et al. (2024b) for details.
@@
For the hard state, we consider the results of fitting with the reflection method to be unreliable due to the assumption adopted in those studies of the disk extending to the ISCO (or to its immediate vicinity). In the soft state, this method can be coupled with the continuum method, making it more reliable, though with the caveats discussed above.
@@ 4.2 Spin-up by accretion
To spin up a BH with an initial mass $M_{0}$ from $a_{*}=0$ to $a_{*}=1$, a mass of $M_{\rm accr}=3M_{0}\arccos\sqrt{2/3}\approx 1.846M_{0}$ needs to be accreted (Bardeen, 1970; Thorne, 1974). It is assumed that the matter accreted carries the specific angular momentum of the ISCO, and the decelerating power of jets (Lowell et al., 2024, 2025) is neglected. However, the BH mass grows by a lower amount, $M_{1}-M_{0}=(6^{1/2}-1)M_{0}\approx 1.449M_{0}$, due to the rest energy being converted to the binding energy. Further accretion does not increase the spin. In terms of the final BH mass, the accreted fraction needed to spin up to $a_{*}=1$ is $M_{\rm accr}/M_{1}=\sqrt{2/3}\arccos\sqrt{2/3}\approx 0.754$, while $M_{0}/M_{1}=6^{-1/2}\approx 0.408$. The actual maximum spin is lower than 1 due to the radiation swallowed by the BH, $a_{*{\rm max}}\approx 0.998$ (Thorne, 1974). For $M_{1}/M_{0}\leq 6^{1/2}$, a null initial spin and an arbitrary final spin, $a_{*}$, we have (Bardeen, 1970; Thorne, 1974)
@@ 4.2.1 BH HXMBs
Some BH HMXBs could be progenitors of binary BH mergers (Gallegos-Garcia et al., 2022), so the need to explain the discrepancy is most immediate for them. In the studies of Cyg X-1, LMC X-1, and M33 X-7 of Zhao et al. (2021b); Gou et al. (2009); Liu et al. (2008), respectively, the BH spin is determined using the continuum method (Section 3.1.2). Those authors assumed that the high spins they measured were natal, which disagrees with the current views (Section 2; Fuller and Ma 2019). The revised studies of those binaries of Belczyński et al. (2024) and Zdziarski et al. (2024a, b) confirmed that their spins are still high (though somewhat lower than before) when fitted by the standard disk models. However, the short lifetimes of HMXBs prevent Eddington-limited accretion from substantially increasing the BH mass (and thus the spin). The Eddington time, i.e., the e-folding time for the mass increase, is independent of the BH mass,
@@
Summarizing this section, the main open questions for the claimed high spins of BH HMXBs are whether they could be natal and, if not, at which stage of the binary evolution the BH spins up via hypercritical accretion, and why those options do not work for BBHs. On the other hand, the actual HMXB spins could be low in alternative disk models, in particular those including warm coronae.
@@ 4.2.2 BH LXMBs
As stated in Section 3.5, the published values of the spins of BH LMXBs span the range from $\sim$0.1 to >>0.99, with the mean value of $\langle a_{*}\rangle>0.7$. There are also some published claims of negative spins, for which the evidence appears relatively uncertain, with a detailed discussion provided in Zdziarski et al. (2025). Assuming their BHs were born with low spin, they were spun up by accretion from the donor. In the first study of that process, King and Kolb (1999) found that the Eddington-limited mass transfer rates and the theoretical lengths of the mass transfer rates in BH XRBs implied that the accretion could not significantly change the BH spins. This finding was revised in Podsiadlowski et al. (2003), who found a spin-up to $a_{*}\lesssim 0.9$ to be possible, provided the masses of the donors at the beginning of the mass transfer onto the BH were high, up to $M_{\rm d}\sim 10{M}_{\odot}$. The difference concerning King and Kolb (1999) was explained by taking into account both an increase of the Eddington rate at low spins (due to the low accretion efficiency) and an increase of the donor evolutionary lifetime as its mass decreases due to the mass transfer. Then, Fragos and McClintock (2015) studied the same problem, but they assumed fully conservative accretion (i.e., not Eddington-limited). With this assumption, it was possible to obtain the BHs with spins up to $a_{*}\approx 1$ and to reproduce all of the 16 Galactic BH XRBs with measured and estimated spins known at the time of their work. We note that the presence of conservative accretion is highly uncertain for phases with supercritical mass transfer, as discussed in Section 4.2.1. Fragos and McClintock (2015) also considered the case of Eddington-limited rates, for which their results were consistent with those of Podsiadlowski et al. (2003). Thus, they were unable to explain $a_{*}\gtrsim 0.9$, whereas $a_{*}\simeq 0.99$ has been claimed for some LMXBs.
@@
The initial donor masses in the models of Podsiadlowski et al. (2003) and Fragos and McClintock (2015) are often much higher than $\sim 1{M}_{\odot}$ typically observed in BH LMXBs. Those intermediate/high-mass donors will quickly lose most of their initial mass through mass loss, as the evolutionary time increases rapidly with decreasing mass. Thus, the probability of seeing an LXMB in the initial phases of mass transfer is very low. On the other hand, it is not clear why low initial donor masses do not strongly dominate the BH LMXB population, given the steepness of the initial stellar mass function, $\propto M^{-2.7}$ above $1{M}_{\odot}$ (Kroupa, 2002), which issue was not discussed in Fragos and McClintock (2015).
@@ 4.3 BBHs and BH HMXBs as two different binary populations
In the case of the known BH HMXBs, the future evolution of Cyg X-1 will certainly lead to the formation of a BBH system, but the BHs will not merge in the Hubble time if the accretion is Eddington-limited (Ramachandran et al., 2025). For fully conservative accretion, the system will merge within $\gtrsim$5 Gyr (Ramachandran et al., 2025). M33 X-7 will, most likely, undergo a common-envelope stage during which the donor and the BH will merge; thus, it will not become a BBH (Ramachandran et al., 2022). While the nature of the compact object in Cyg X-3 is not certain, it is most likely a BH (Antokhin et al., 2022). However, its masses remain relatively uncertain. If the mass of the WR donor is $\gtrsim 13{M}_{\odot}$, it will become a BBH within less than 1 Myr, and the BHs will most likely merge (Belczyński et al., 2013). However, as discussed in that paper, its final fate is relatively uncertain.

### v3

@@ 3.3 Jet power and spins
Transient ejecta achieve the highest power among jets launched by BH XRBs during hard-to-soft state transitions, with a prominent example of MAXI J1348–630 (Carotenuto et al., 2021). The jet power has been estimated based on the kinetic energy obtained from modeling the propagation of those jets (observed in radio) up to distances up to $\sim$1 pc through the surrounding media, together with the jet launching times assumed to equal the durations of major radio flares accompanying the ejecta. Typically, the major flares are usually well time-resolved, and last $\lesssim$1 day (Carotenuto et al., 2021; Cooper et al., 2025). The resulting values of the power are comparable to the maximum power of Equation (8) at $a_{*}\sim 1$ in several cases (Carotenuto et al., 2024; Cooper et al., 2025). The vast propagation distances require the presence of surrounding cavities with the densities $\ll 1$ cm-3 (Heinz, 2002; Sikora and Zdziarski, 2023). Notably, such powerful ejections have been observed only among LMXBs. Thus, some of them may be required to have high spins. On the other hand, Fender and Motta (2025) have compared the measured ejecta velocities with the published values of the spin obtained using the reflection, continuum, and QPO methods, and found no correlations. This shows that either the ejecta power is at most weakly dependent on the spin, or the published spin values are unreliable (which we advocate here), or both.
@@ 3.4 X-ray polarization and spins
An independent method for measuring the BH spin is through X-ray polarization. The image of the innermost regions of the accretion disk is distorted by general and special relativity effects, causing different parts of the disk – whose emission dominates at different energies – to have varying orientations. This effect is expected to produce a gradual change of PA as a function of energy and progressive decrease of PD (Dovčiak et al., 2004a; Loktev et al., 2022, 2024). The rate of PA rotation and the suppression of PD are sensitive to BH spin, making them valuable diagnostic tools. However, X-ray polarization measurements of soft-state sources show neither statistically significant evidence of PA dependence on energy, nor do they indicate a decrease of PD with increasing energy (Svoboda et al., 2024; Marra et al., 2024; Ratheesh et al., 2024; Steiner et al., 2024).
The importance of returning radiation was proposed to explain the observed increase of PD with energy and the nearly constant PA in two sources, 4U 1957+115 and Cyg X-1 (Marra et al., 2024; Steiner et al., 2024). However, satisfying both criteria for the latter required an assumption of an extreme albedo of unity. For the former, the albedo was found to be not constrained. Moreover, no simultaneous fits of both the spectral and polarimetric signatures were done for the latter: either spectral modeling was performed separately from polarimetric modeling, or the parameters that provided a good fit for the polarization data failed to align with the spectral constraints.
@@ 3.5 Published values of the BH spins of XRBs
We know three BH HMXBs with measured spins: Cyg X-1, LMC X-1, and M33 X-7. Their spin values published before 2024 were all high: >>0.9985 (Zhao et al., 2021b; Miller-Jones et al., 2021), $0.92^{+0.05}_{-0.07}$ (Gou et al., 2009), and $0.84\pm 0.05$ (Liu et al., 2008), respectively. Those results were obtained using the disk continuum method in the soft state, specifically with kerrbb2. A new value for Cyg X-1 was found in the spectral fit with kerrbb by Steiner et al. (2024), $a_{*}=0.99964^{+0.00003}_{-0.00007}$. Although those errors do not include systematic errors and uncertainties in the mass, distance, and inclination, very high values of the spin of Cyg X-1 were also reported using the previous determination of the mass and distance. In particular, Gou et al. (2014) found $a_{*}>0.983$ using the values of $M=14.8{M}_{\odot}$ and $D=1.86$ kpc (Orosz et al., 2011).
Similarly, the published spin values for BH LMXBs are high on average, with $a_{*}\gtrsim 0.9$ in many cases and the mean value of $\langle a_{*}\rangle>0.7$; see the compilations in Reynolds (2021) and Fishbach and Kalogera (2022). A recent paper by Draghis et al. (2024) used the reflection method for a large sample of BH XRBs and claimed that the spins obtained in their work are even higher for most sources than those found previously. Namely, about 86% of their sample had spins consistent with $a_{*}\geq 0.95$, 94% were consistent with $a_{*}\geq 0.9$, and 100% of their spin values were consistent with $a_{*}>0.7$. However, they fitted models with the incident spectrum being either an e-folded power law or a single thermal Comptonization spectrum. As we argued in Section 3.1.1, the actual incident spectra are significantly more complex. In particular, spectral fits in the hard state with two Comptonization components typically yield truncated disks (e.g., Basak et al. 2017; Zdziarski et al. 2021a, b; Chand et al. 2024; Sahu et al. 2025), which then do not allow a spin measurement. On the other hand, Draghis et al. (2024) assumed that the disk extends down to the ISCO in all cases, in both the hard and soft states. Taking their results at face value, even larger spins would be implied, with the obtained values being lower limits.
@@
Then, the spins have been measured to be low in some binaries while they showed powerful jets, apparently requiring high spins to power them (Equation 8 in Section 3.3). A prominent case is that of MAXI J1820+070. Using the continuum method, Guan et al. (2021) and Zhao et al. (2021a) obtained $a_{*}=0.2^{+0.2}_{-0.3}$ and $a_{*}=0.14\pm 0.09$, respectively, while Fabian et al. (2020) argued for $a_{*}<0$ based on the possible presence of emission from the plunging region below the ISCO (see Section 3.1.2). On the other hand, this system showed a powerful transient jet (Bright et al., 2020; Carotenuto et al., 2024), which requires $a_{*}\sim 1$ (Zdziarski and Heinz, 2024).
@@ 4.1 Revised spin measurements and improved disk models
Figure 6: A soft-state state of Cyg X-1 observed by NICER, NuSTAR and INTEGRAL (based on Zdziarski et al. 2024b). We show the unfolded spectra and data-to-model ratios in the top and bottom panels. In (a), the spectrum is fitted by the disk model kerrbb and Comptonization, yielding $a_{*}=0.87^{+0.04}_{-0.03}$. In (b), the spectrum is fitted with a similar model but including a warm corona, and $a_{*}=0.00^{+0.07}$. The total model spectra and the unabsorbed ones are shown by the solid blue and blue curves, respectively. The unabsorbed disk, scattered, and reflected components are shown by the magenta, green, and red curves, respectively. In (b), the brown curve represents the underlying disk spectrum before it undergoes Comptonization in the top layer. See Zdziarski et al. (2024b) for details.
@@
For the hard state, we consider the results of fitting with the reflection method to be unreliable due to the assumption adopted in those studies of the incident spectrum to be a single power law, while allowing more complex spectra (e.g., two Comptonization components) often yield truncated disks (which do not allow a spin measurement). In the soft state, this method can be coupled with the continuum method, making it more reliable, though with the caveats discussed above.
@@ 4.2 Spin-up by accretion
To spin up a BH with an initial mass $M_{0}$ from $a_{*}=0$ to $a_{*}=1$, a mass of $M_{\rm accr}=3M_{0}\arccos\sqrt{2/3}\approx 1.846M_{0}$ needs to be accreted (Bardeen, 1970; Thorne, 1974). It is assumed that the matter accreted carries the specific angular momentum of the ISCO, and the decelerating power of jets (Lowell et al., 2024, 2025) is neglected. However, the BH mass grows by a lower amount, $M_{1}-M_{0}=(6^{1/2}-1)M_{0}\approx 1.449M_{0}$, due to the rest energy being converted to the binding energy. Any additional accretion does not increase the spin. In terms of the final BH mass, the accreted fraction needed to spin up to $a_{*}=1$ is $M_{\rm accr}/M_{1}=\sqrt{2/3}\arccos\sqrt{2/3}\approx 0.754$, while $M_{0}/M_{1}=6^{-1/2}\approx 0.408$. The actual maximum spin is lower than 1 due to the radiation swallowed by the BH, $a_{*{\rm max}}\approx 0.998$ (Thorne, 1974). For $M_{1}/M_{0}\leq 6^{1/2}$, a null initial spin and an arbitrary final spin, $a_{*}$, we have (Bardeen, 1970; Thorne, 1974)
@@ 4.2.1 BH HXMBs
Some BH HMXBs could be progenitors of binary BH mergers, so the need to explain the discrepancy is most immediate for them. In the studies of Cyg X-1, LMC X-1, and M33 X-7 of Zhao et al. (2021b); Gou et al. (2009); Liu et al. (2008), respectively, the BH spin is determined using the continuum method (Section 3.1.2). Those authors assumed that the high spins they measured were natal, which disagrees with the current views (Section 2; Fuller and Ma 2019). The revised studies of those binaries of Belczyński et al. (2024) and Zdziarski et al. (2024a, b) confirmed that their spins are still high (though somewhat lower than before) when fitted by the standard disk models. However, the short lifetimes of HMXBs prevent Eddington-limited accretion from substantially increasing the BH mass (and thus the spin). The Eddington time, i.e., the e-folding time for the mass increase, is independent of the BH mass,
@@
Summarizing this section, the main open questions for the claimed high spins of BH HMXBs are whether they could be natal and, if not, at which stage of the binary evolution the BH spins up via hypercritical accretion, and why those scenarios do not appear in BBHs. On the other hand, the actual HMXB spins could be low when fitted by alternative disk models, in particular those including warm coronae.
@@ 4.2.2 BH LXMBs
As stated in Section 3.5, the published values of the spins of BH LMXBs span the range from $\sim$0.1 to >>0.99, with the mean value of $\langle a_{*}\rangle>0.7$. There are also some published claims of negative spins, for which the evidence appears relatively uncertain, with a detailed discussion provided in Zdziarski et al. (2025). Assuming their BHs were born with low spin, they were spun up by accretion from the donor. In the first study of that process, King and Kolb (1999) found that the Eddington-limited mass transfer rates and the theoretical lengths of the mass transfer rates in BH XRBs implied that the accretion could not significantly change the BH spins. This finding was revised in Podsiadlowski et al. (2003), who found a spin-up to $a_{*}\lesssim 0.9$ to be possible, provided the masses of the donors at the beginning of the mass transfer onto the BH were high, up to $M_{\rm 2,0}\sim 10{M}_{\odot}$. The difference concerning King and Kolb (1999) was explained by taking into account both an increase of the Eddington rate at low spins (due to the low accretion efficiency) and an increase of the donor evolutionary lifetime as its mass decreases due to the mass transfer. Then, Fragos and McClintock (2015) studied the same problem, but they assumed fully conservative accretion (i.e., not Eddington-limited). With this assumption, it was possible to obtain BHs with spins up to $a_{*}\approx 1$ and to reproduce all of the 16 Galactic BH XRBs with measured and estimated spins known at the time of their work. We note that the presence of conservative accretion is highly uncertain for phases with supercritical mass transfer, as discussed in Section 4.2.1. Fragos and McClintock (2015) also considered the case of Eddington-limited rates, for which their results were consistent with those of Podsiadlowski et al. (2003). Thus, they were unable to explain $a_{*}\gtrsim 0.9$, whereas $a_{*}\simeq 0.99$ has been claimed for some LMXBs.
@@
The initial donor masses in the models of Podsiadlowski et al. (2003) and Fragos and McClintock (2015) are often much higher than $\sim 1{M}_{\odot}$ typically observed in BH LMXBs. Those intermediate/high-mass donors will quickly lose most of their initial mass through mass loss, as the evolutionary time increases rapidly with decreasing mass. Thus, the probability of seeing an LXMB in the initial phases of mass transfer is very low. The initial mass distribution of companions in stellar binaries spans a wide range (Sana et al., 2012), although it may exhibit some dependencies on the initial orbital period (Moe and Di Stefano, 2017). The initial orbital periods of the progenitors of LMXBs remain, however, poorly constrained due to the absence of a fully convincing mass-transfer scenario that can robustly reproduce the observed orbital separations of LMXBs within the framework of classical common-envelope evolution (e.g., Kalogera, 1999; Podsiadlowski et al., 2003; Wiktorowicz et al., 2014; Wang et al., 2016). An additional potentially important ingredient is the high fraction of triple systems among massive stars (Moe and Di Stefano, 2017). The presence of a tertiary companion may influence the early evolution of the inner binary through secular dynamical effects, such as the eccentric Kozai–Lidov mechanism (Naoz et al., 2016), thereby providing alternative pathways toward the formation of BH LMXBs.
Binaries with low-mass companions ($M_{2}\sim 1{M}_{\odot}$) may be more easily disrupted during BH formation, for example, due to the Blaauw kick associated with sudden mass loss (Blaauw, 1961). This effect could reduce the survival probability of such systems and suppress their contribution to the observed BH LMXB population, thus favoring intermediate-mass companions as the progenitors of present-day BH LMXBs. On the other hand, another population of systems hosting BH candidates–including all three detached Gaia BH systems announced so far (El-Badry et al., 2023b; Chakrabarti et al., 2023; El-Badry et al., 2023a; Gaia Collaboration et al., 2024)–contains companions with masses of the order of $\sim 1{M}_{\odot}$. These systems are found in the Galactic field and show no evidence for triple or higher-order multiplicity; their properties also pose a significant challenge for formation through conventional mass-transfer scenarios (Nagarajan et al., 2025; Olejak et al., 2025). Consequently, both observations and evolutionary models currently provide only weak constraints on the initial masses of the lower-mass companions in LMXB progenitors.
@@ 4.3 BBHs and BH HMXBs as two different binary populations
In the case of the known BH HMXBs, the future evolution of Cyg X-1 will certainly lead to the formation of a BBH system, but the BHs will not merge in the Hubble time if the accretion is Eddington-limited. Only for fully conservative accretion, the system will merge within $\gtrsim$5 Gyr (Ramachandran et al., 2025). M33 X-7 will, most likely, undergo a common-envelope stage during which the donor and the BH will merge; thus, it will not become a BBH (Ramachandran et al., 2022). While the nature of the compact object in Cyg X-3 is not certain, it is most likely a BH (Antokhin et al., 2022). However, its masses remain relatively uncertain. If the mass of the WR donor is $\gtrsim 13{M}_{\odot}$, it will become a BBH within less than 1 Myr, and the BHs will most likely merge (Belczyński et al., 2013).
The BHs in mergers of field binaries also have undergone evolution in mass and spin, and in this respect are similar to XRBs. The crucial issue, which we stress in the paper, is that the GW results strongly imply that the natal spins are low. To explain the spins measured in HMXBs, the natal spin has to be very high, in contrast to similar BHs undergoing mergers. No viable explanation has been proposed yet. For LMXBs, either the natal spins are high, which again is unexplained, or the donor masses at the onset of the RLO were high, which remains to be shown, and which we discuss in detail.
It has been proposed that failed supernova explosions can spin up BHs through the accretion of the remaining hydrogen envelope or fallback material. Antoni and Quataert (2022) considered accretion from the hydrogen envelope of a red supergiant progenitor and found that this mechanism does not produce spins higher than $a\simeq 0.8$. On the other hand, Batta et al. (2017), focusing on HMXBs, suggest that spins exceeding $a\gtrsim 0.8$ can be reproduced if a substantial amount of fallback material is generated during the failed explosion, and if its interaction with the companion star increases its angular momentum content. Once this material is eventually accreted, it may lead to a rapidly spinning BH. We argue, however, that the failed supernova spin-up mechanism is unlikely to resolve the apparent spin tension between the GW and HMXB populations, and may be challenged for several reasons.
In particular, at least a fraction of BBH merger events—especially those involving very massive BHs (Fryer et al., 2012)—are also expected to form through failed supernovae or direct collapse. The associated low natal kicks in such explosions would help binaries avoid disruption and merge within a Hubble time. Yet current GW observations do not show evidence for high spins, and certainly not for spins approaching $a\sim 0.8$. Moreover, if failed supernovae were the dominant pathway to producing highly spinning BHs, one would expect a correlation between BH mass and final spin above some threshold for direct-collapse formation (Fryer et al., 2012). While the most massive BHs in the current BBH sample (with masses $\gtrsim 30\,{M}_{\odot}$) may show hints of moderately elevated spins (The LIGO Scientific Collaboration et al., 2025c), these masses are significantly higher than those inferred for X-ray binaries. BHs detected via GWs with masses comparable to typical XRB BHs ($\lesssim 20\,{M}_{\odot}$) instead remain consistent with $a_{\rm eff}\sim 0$ and low individual spins.
Finally, as discussed above, BH progenitors in both populations undergo mass-transfer phases that likely remove a substantial fraction of the hydrogen envelope. In addition, XRBs are frequently found in relatively high-metallicity environments, where stellar winds can remove up to $\sim 80\%$ of the progenitor’s initial mass (Belczynski et al., 2010), carrying away a significant amount of angular momentum. If the failed supernova spin-up scenario were operating efficiently, one would naively expect GW sources—often forming in lower-metallicity environments, experiencing weaker stellar winds (and thus more likely retaining part of their hydrogen envelopes)—to host more rapidly spinning BHs than those observed in XRBs. This expectation is contrary to current observations.